OpenStudy (dan815):

why is it x instead of 1x

OpenStudy (dan815):

nvm i figured it out

OpenStudy (mrnood):

not sure what this refers to - is there an equation or diagram to post? However - x and 1x are the same in normal notation we write 2x or 3x etc to mean 2 times x but 1 times x is just x so we normally missout the 1 for clarity

OpenStudy (dan815):

its because 2x represents x+x and 1x represents x so 1x is the only one that can simplify. 2x and all other x are not simpler except 0x which is nothing.

OpenStudy (cal.lavender):

mew

OpenStudy (dan815):

2x and all the other kx* are not simpler except 0x which is nothing.

OpenStudy (cal.lavender):

mew

OpenStudy (cal.lavender):

@dan815 's a kitty

ganeshie8 (ganeshie8):

reverse addition, multplication notation and see how the distribution law looks

ganeshie8 (ganeshie8):

multiplication : + addition : nothing

ganeshie8 (ganeshie8):

looks we can't express it w/o using parenthesis \[a+bc = (a+b)(a+c)\]

OpenStudy (shadowlegendx):

Dan r u dumbo?

OpenStudy (shadowlegendx):

I learned this in middle school

OpenStudy (anonymous):

i thought we was to lazy to put a 1

OpenStudy (dan815):

no think about it!! its just an x

OpenStudy (dan815):

2x is a representation of x+x 1x is a representation of x why would you write 1x for x when x is what it's already referring to is being used in the definition of it.

OpenStudy (anonymous):

oh cool

OpenStudy (phi):

1x means 1*x and that "simplifies" to x the same way 1*2 = 2 or 1*6= 6 you can always multiply by 1, but usually you don't bother to do that. But sometimes, it makes things clearer For example (x+ 2yx) we can "factor out an x" but the process is less mysterious if we show that x can be written 1*x and we have (1*x + 2y*x) and (1+2y)*x is the factored form.